A sphere has a diameter of 12m. Find its volume using Pie=3.14 as an approximation
Answer by checkley71(8403)

V=4/3PIR^3
V=4/3*3.14*6^3
V=4.1867*216
V=904.33 M^3 IS THE VOLUME.